# Results and Lessons Learned from Autonomous Driving Transportation Services in Airfield, Crowded Indoor, and Urban Environments

### 02 Mar 2024 v1

@@ III-C Urban Environment
As shown in Figs. 1c and 2c, We have been operating various autonomous driving services, including last-mile delivery, health care, and infra monitoring, in urban environments. Here we describe the novel algorithms we proposed for product-level autonomous shuttle and taxi services in a 31.8 km route in Jinhae, South Korea. The shuttle operates between 19 stops and the taxi operates at 23 stops. These services require real-time and high-accuracy perception performance and safe and interactive driving behaviors. We conducted a performance evaluation with a total distance of 110.3 km, without any disengagement. Table III presents the autonomous shuttle test statistics in Jinhae.
@@ III-C1 Imagination-augmented Hierarchical Reinforcement Learning for Safe and Interactive Motion Planning
Many previous works demonstrated that rule-based motion planning algorithms can perform safety-aware behaviors following traffic rules in urban environments [9, 10]. However, as these algorithms are pre-scripted for specific scenarios and cannot understand interactions with human drivers, their behaviors are often much inferior to the behaviors of expert drivers. While learning-based motion planning algorithms are a promising approach for training autonomous vehicles to understand interactions with human drivers, their behaviors do not necessarily guarantee safety.
@@
where $\sigma$ is a softmax function, $Q$, $K$ and $V$ are the query, key, and value matrices, respectively, $s_{t}$ is a state, $\zeta_{t}^{ego}$ is the imagined behavior of the ego vehicle, and $\eta$ is a random seed vector. The state consists of imagined behaviors of the ego vehicle and surrounding vehicles $s_{t}=\{\zeta_{t}^{ego},\zeta_{t}^{1},\zeta_{t}^{2},\dots,\zeta_{t}^{n}\}$. Note that the key and value matrices are linear transformations of $s_{t}$, and the query matrix is linear transformations of $\zeta_{t}^{ego}$. The experimental results showed that IAHRL achieved a success rate of over 95 % in various unsignalized intersections and lane-change tasks, significantly outperforming baselines. We also observed that IAHRL enables an autonomous vehicle to give its attention to surrounding vehicles in a manner akin to human drivers.
@@ III-C2 Real-time and Accurate Multi-sensor Object Detection
Object detection algorithms for autonomous vehicles must ensure real-time and high-accuracy with limited computation capacity. Cross modal transformer (CMT) [26], a state-of-the-art sensor fusion algorithm, showed a high performance by utilizing transformers to fuse the features of LiDAR and multiple cameras. Nevertheless, it has limitations, such as dependence on camera intrinsic parameters and a high computational power requirement.
We developed a multi-sensor object detection (MSOD) algorithm that achieves high performance while satisfying real-time constraints. MSOD extracts bird’s-eye view features from LiDAR data and employs only the corresponding image features as the transformer input. Thus, it can more efficiently operate than the CMT, which uses all LiDAR and camera features as the input. We designed the model to leverage the advantages of each sensor by utilizing LiDAR features for depth information and camera features for semantic information, ensuring high accuracy. Another distinctive feature of MSOD is its ability to operate with LiDAR alone even under camera failure.
Open datasets, such as KITTI and nuScene, differ from the Korean urban environment; hence, we generated a dataset using our autonomous driving platform (Fig. 2) to accurately evaluate the performance. The evaluation metrics were nuScene detection score (NDS) and mean average precision (mAP). We compared inference speed tested on RTX3090 GPU using Python. Fig. 5 and Tabel IV present the qualitative and quantitative results, respectively. The experimental results indicate that our algorithm, while maintaining accuracy comparable to CMT, is nearly twice as fast. MSOD is converted to TensorRT and integrated into vehicles, operating at a frequency of 28 Hz.
@@ IV-A1 Jetblast Detection
Jetblast is an intangible rapid airflow from an aircraft’s engines and typically occurs when it prepares for takeoff or right after landing. Autonomous tow tractors should be able to avoid the jetblast, as it can cause significant damage to them. However, most previous works on autonomous driving have focused on detecting objects such as vehicles and pedestrians that are common in urban driving environments. We observed that their algorithms are inefficient in detecting the jetblast due to its unique features. To overcome this challenge, we are currently developing a learning-based jetblast detection algorithm using thermal cameras (Fig. 7). We use a YOLO-based 2D detection network utilizing the temperature value of the thermal camera as a feature. Moreover, we believe that incorporating sound localization through audio detection can further enhance the jetblast detection accuracy [27].
@@ IV-A2 Robust and Adaptive Control for Variable Tow Tractor Dynamics
The tow tractor’s payloads can change as shown in Fig. 7, and it significantly affects the dynamics of tow tractors. To ensure the efficiency and safety of autonomous tow tractors, a robust control algorithm that can adapt to changing dynamics is essential. Most previous control algorithms for tow tractors handle this challenge by assuming a prior knowledge of the tractor’s configuration. However, the tractor’s configuration varies frequently and cannot be predicted in advance. Meta-learning that facilitates effective and continuous online adaptation is a promising approach for handling the changing dynamics problem [28]. Nagabandi et al. [29] introduced a model-based meta-learning algorithm that can continuously adapt a model to recent transitions in dynamic environments. We believe that integrating the benefits of meta-learning into our controller can enable tow tractors to adapt to changing dynamics due to payloads.
@@ IV-B1 Human-Robot Interaction for Social Navigation
In general, indoor robots require navigation algorithms to avoid interrupting people and move efficiently [30]. AirRide infers efficient and adaptive paths without interrupting surrounding people by predicting their behaviors in advance. However, AirRide was occasionally frozen when people standing in line for boarding blocked whole passages, which is common at arrival and departure halls. While AirRide is programmed to stop and then request yielding through voice guidance when it encounters these scenarios, we observed that our voice guidance was not sufficient to enable AirRide to encourage people to cooperate. We believe that investigating socially compliant driving is a good starting point for handling our limitation [31]. We thus will analyze how existing socially compliant driving algorithms work in the challenging scenario at the arrival and departure halls.
IV-B2 Domain Gap Between the Simulator and the Real World
A simulator is essential for deploying autonomous vehicles in real-world environments. There are numerous simulators representing urban environments at high fidelity [32]. However, available simulators for crowded environments have many shortcomings [33]. One of the main limitations is that simulators cannot accurately model the diverse behaviors of people. For instance, at airports, people may pull their luggage, push carts with luggage, use wheelchairs, or move with young children in various ways. Simulators also do not support modeling the behavior patterns resulting from people’s interest in autonomous vehicles. Camara et al. [34] introduced realistic modeling of people’s behavior, which can significantly improve the simulator performance. Minimizing such domain gap between simulators and real-world environments could be of great help in autonomous driving research in crowded indoor environments.
@@ IV-C1 Human Intervention for Training Autonomous Vehicles
While IAHRL introduced in Section III.C.1) enables autonomous vehicles to learn safe and interactive driving behaviors in urban environments, a lot of additional challenges should be addressed to truly facilitate the training of autonomous vehicles in real-world environments. One challenge that has not received much attention is the substantial human intervention and supervision required for training autonomous vehicles. Humans should continuously supervise autonomous vehicles during training to intervene before they enter unsafe states. Furthermore, after each episode, humans should reset environments by manually driving autonomous vehicles to the initial states for the next episode. Autonomous RL (ARL), which learns not only how to solve a task but also how to reset an environment, is a promising approach to minimizing human intervention and supervision. Several recent works introduced impressive ARL algorithms that can substantially reduce human intervention in diverse tasks [35]. However, applying these algorithms to train an autonomous vehicle in real-world environments is infeasible, as they cannot guarantee safety-aware reset behaviors required in autonomous driving tasks. We believe that developing ARL algorithms that can safely reset an environment is crucial for realizing autonomous vehicles in the real world.
@@ IV-C3 Time Gap Between Environment Change and Deploying Updated Maps
A precise high-definition (HD) map is used in many components of autonomous driving systems, such as map matching-based localization, trajectory prediction, and global path planning algorithms. However, continuously and autonomously updating the HD maps to represent environmental changes is not straightforward in the real world. The outdated HD maps can lead to significantly poor autonomous driving performance. To overcome this challenge, Kim et al. [38] introduced a crowd-sourcing algorithm that leverages a large number of vehicle fleets with low-cost sensors to continuously update the map in real-time. However, even with the fast map update through crowd-sourcing, there is a time gap between environment change and deploying updated maps to autonomous vehicles. A fail-operational system, maintaining continuous operation during system failures, is one of the promising solutions that can handle this time gap, but it has yet to receive much attention. We believe that switching to a perception-based autonomous driving mode, which relies solely on current perception results without HD maps, when environment change is detected, is an appropriate fail-operational system.
@@ V Conclusion
We share our experimental results and lessons obtained from autonomous driving services in airfield, crowded indoor, and urban environments. Each environment has its own unique challenging problems. We introduce general and novel algorithms for addressing several problems and provide an overview of promising research directions for the remaining problems that have not received much attention but must be addressed. Exploring these research directions is crucial for realizing the massive potential of autonomous vehicles. We hope that our work will be helpful to researchers who would like to deploy autonomous vehicles in various real-world environments.

### 20 Mar 2024 v2

@@ III-C Urban Environment
As shown in Figs. 1c and 2c, we have been operating various autonomous driving services, including last-mile delivery, health care, and infra monitoring, in urban environments. Here we describe the novel algorithms we proposed for product-level autonomous shuttle and taxi services in a 31.8 km route in Jinhae, South Korea. The shuttle operates between 19 stops and the taxi operates at 23 stops. These services require real-time and high-accuracy perception performance, and safe and interactive driving behaviors. We conducted a performance evaluation with a total distance of 110.3 km, without any disengagement. Table III presents the autonomous shuttle test statistics in Jinhae.
Many previous works demonstrated that rule-based motion planning algorithms can perform safety-aware behaviors following traffic rules in urban environments [8, 9]. However, as these algorithms are pre-scripted for specific scenarios and cannot understand interactions with human drivers, their behaviors are often much inferior to the behaviors of expert drivers. While learning-based motion planning is a promising approach for training autonomous vehicles to understand interactions with human drivers, it does not necessarily generate safety-aware driving behaviors.
@@ III-C1 Imagination-augmented Hierarchical Reinforcement Learning for Safe and Interactive Motion Planning
where $\sigma$ is a softmax function, $Q$, $K$ and $V$ are the query, key, and value matrices, respectively, $s_{t}$ is a state, $\zeta_{t}^{ego}$ is the imagined behavior of the ego vehicle, and $\eta$ is a random seed vector. The state consists of imagined behaviors of the ego vehicle and surrounding vehicles $s_{t}=\{\zeta_{t}^{ego},\zeta_{t}^{1},\zeta_{t}^{2},\dots,\zeta_{t}^{n}\}$. Note that the key and value matrices are linear transformations of $s_{t}$, and the query matrix is linear transformations of $\zeta_{t}^{ego}$. The experimental results showed that IAHRL achieves success rates of over 95 % in various unsignalized intersections and lane-change tasks, significantly outperforming baselines. We also observed that IAHRL enables an autonomous vehicle to give its attention to surrounding vehicles in a manner akin to human drivers.
@@ III-C2 Real-time and Accurate Multi-sensor Object Detection
Object detection algorithms for autonomous vehicles must ensure real-time and high accuracy with limited computation capacity. Cross modal transformer (CMT) [25], a state-of-the-art sensor fusion algorithm, showed a high performance by utilizing transformers to fuse the features of LiDAR and multiple cameras. Nevertheless, it requires high computational resources, making it difficult to satisfy the real-time constraints of autonomous vehicles.
We developed a multi-sensor object detection (MSOD) algorithm that achieves high performance while satisfying real-time constraints. MSOD extracts bird’s-eye view features from LiDAR data and employs only the corresponding image features as the transformer input. Thus, it can more efficiently operate than the CMT, which uses all LiDAR and camera features as the input. MSOD uses the model to leverage the advantages of each sensor by utilizing LiDAR features for depth information and camera features for semantic information, ensuring high accuracy. Another distinctive feature of MSOD is its ability to operate with LiDAR alone even under camera failure.
Open datasets, such as KITTI and nuScene, differ from the Korean urban environment; hence, we generated a dataset using our autonomous driving platform (Fig. 2) to accurately evaluate the performance. The evaluation metrics were nuScene detection score (NDS) and mean average precision (mAP). We compared inference speed tested on RTX3090 GPU using Python. Fig. 5 and Tabel IV present the qualitative and quantitative results, respectively. The experimental results indicate that our algorithm, while maintaining accuracy comparable to CMT, is nearly twice as fast. We confirmed that MSOD can operate at a frequency of 28 Hz after conversion to TensorRT.
@@ IV-A1 Jetblast Detection
Jetblast is an intangible rapid airflow from an aircraft’s engines and typically occurs when it prepares for takeoff or right after landing. Autonomous tow tractors should be able to avoid the jetblast, as it can cause significant damage to them. Unfortunately, most previous detection algorithms for autonomous driving have focused on detecting objects common in urban environments, such as vehicles and pedestrians. We observed that these algorithms are inefficient in detecting the jetblast due to its unique features. To overcome this challenge, we are currently developing a thermal camera-based jetblast detection algorithm (Fig. 7). Thermal cameras are widely used to detect such intangible objects. We confirmed that jetblast is successfully detected using a YOLO-based 2D detection network with temperature features. Moreover, we believe that incorporating sound localization through audio detection can further enhance the jetblast detection accuracy [26].
@@ IV-A2 Robust and Adaptive Control for Variable Tow Tractor Dynamics
The tow tractor’s payload can change as shown in Fig. 7, and it significantly affects the dynamics of tow tractors. To ensure the efficiency and safety of autonomous tow tractors, a robust control algorithm that can adapt to changing dynamics is essential. Most previous control algorithms for tow tractors handle this challenge by assuming a prior knowledge of the tractor’s configuration. However, the tractor’s configuration varies frequently and cannot be predicted in advance. Meta-learning that facilitates effective and continuous online adaptation is a promising approach for handling the changing dynamics problem [27]. Nagabandi et al. [28] introduced a model-based meta-learning algorithm that can continuously adapt a model to recent transitions in dynamic environments. We believe that integrating the benefits of meta-learning into our controller can enable tow tractors to adapt to changing dynamics due to payload.
@@ IV-B1 Human-Robot Interaction for Social Navigation
In general, navigation algorithms for indoor robots should infer efficient and adaptive paths that do not interrupt people’s behaviors [29]. AirRide can generate such paths by predicting people’s behaviors in advance. However, AirRide was occasionally frozen when people standing in line for boarding blocked whole passages, which is common at arrival and departure halls. While AirRide is programmed to stop and then request yielding through voice guidance when it encounters these scenarios, we observed that our voice guidance was not sufficient to enable AirRide to encourage people to cooperate. We believe that investigating socially compliant driving is a good starting point for handling our limitation [30]. We thus will analyze how existing socially compliant driving algorithms work in the challenging scenario at the arrival and departure halls.
IV-B2 Domain Gap Between Simulator and the Real World
A simulator is essential for evaluating autonomous vehicles before deploying them in real-world environments. While numerous simulators have been developed for urban environments and have achieved high fidelity [31], those designed for crowded indoor environments still have many shortcomings [32]. One of the main limitations is the inaccurate modeling of people’s diverse behaviors, such as pulling their luggage and pushing carts at airports. Specifically, behaviors resulting from people’s interest in autonomous vehicles have been overlooked in most simulators. Camara et al. [33] introduced realistic modeling of people’s behavior, ranging from simple individual prediction to game-theoretic models of interactions between people and autonomous vehicles. This realistic modeling can significantly reduce the domain gap between simulators and real-world environments, which is crucial for advancing autonomous driving research in crowded indoor environments.
@@ IV-C1 Human Intervention for Training Autonomous Vehicles
While IAHRL introduced in Section III.C.1) enables autonomous vehicles to learn safe and interactive driving behaviors in urban environments, a lot of additional challenges should be addressed to truly facilitate the training of autonomous vehicles in real-world environments. One challenge that has not received much attention is the substantial human intervention and supervision required for training autonomous vehicles. Humans should continuously supervise autonomous vehicles during training to intervene before they enter unsafe states. Furthermore, after each episode, humans should reset an environment by manually driving autonomous vehicles to the initial states for the next episode. Autonomous RL (ARL), which learns not only how to solve a task but also how to reset an environment, is a promising approach to minimizing human intervention and supervision. Several recent works introduced impressive ARL algorithms that can substantially reduce human intervention in diverse tasks [34]. However, applying these algorithms to train an autonomous vehicle in real-world environments is infeasible, as they cannot guarantee safety-aware reset behaviors required in autonomous driving tasks. We believe that developing ARL algorithms that can safely reset an environment is crucial for realizing autonomous vehicles in the real world.
@@ IV-C3 Time Gap Between Environment Change and Deploying Updated Maps
High-definition (HD) maps are used in various components of autonomous driving systems, including map matching-based localization, trajectory prediction, and global path planning algorithms. Outdated HD maps can significantly degrade autonomous driving performance. However, the continuous and autonomous updating of HD maps to reflect environmental changes is not straightforward in the real world. To overcome this challenge, Kim et al. [37] introduced a crowd-sourcing algorithm that leverages a large number of vehicle fleets with low-cost sensors to continuously update the maps in real-time. However, even with the fast map update through crowd-sourcing, there is a time gap between environment change and deploying updated maps to autonomous vehicles. A fail-operational system, maintaining continuous operation during system failures, is one of the promising solutions that can handle this time gap, but it has yet to receive much attention. We believe that switching to a perception-based autonomous driving mode, which relies solely on current perception results without HD maps, when environment change is detected, is an appropriate fail-operational system.
@@ V Conclusion
We share our experimental results and lessons obtained from autonomous driving services in airfield, crowded indoor, and urban environments. Each environment has its own unique challenging problems and we introduce general and novel algorithms for addressing them. We also provide an overview of promising research directions for the remaining problems that have not received much attention but must be addressed. Exploring these research directions is crucial for realizing the massive potential of autonomous vehicles. We hope that our work will be helpful to researchers who would like to deploy autonomous vehicles in various real-world environments.